Hi all, can anyone please show me how to do the following question?
Determine whetherher the following line [(x-3)/5] = [(y+1/-2] = [(x-3)/4] contains the point (3, -1, 1).
Steps and notes on how to solve would be really helpful.
Thanks, -nick

Answer by Alan3354(69443)

Determine whetherher the following line [(x-3)/5] = [(y+1/-2] = [(x-3)/4] contains the point (3, -1, 1).
----------------
Sub x, y and z (I assume you meant z in the last term).
[(x-3)/5] = [(y+1/-2] = [(x-3)/4]
0/5 = 0/-2 = -2/4
It's not equal, it's not on the line.